### v2

@@ 5 Preliminary Results and Discussion
The dilemma decision accuracy reached 100%, meaning that the LLM correctly interpreted the natural language expression of the agent’s strategy to choose an action. Similarly, punishment accuracy was highly reliable across all strategies and conditions, indicating that agents correctly identified and punished defectors and non-punishers as per their strategies. Furthermore, this suggests that the influence of agent lifestyles was secondary. The LLM agents aligned their decisions with the behavioural strategy, showing no bias from additional factors such as lifestyles.
@@
The results of the first experimental setup show that without explicit punishment values, the defecting agents (agents with R1) have overtaken the population by the second iteration, as shown in Figure 9(a). The other cooperative strategies (M, P, E) have been replaced with R1 strategies due to lower utilities compared to their R1 counterparts. This indicates that the LLM may inherently choose punishment costs (p and k) that are not sufficient to override the R1 strategy when not specifying p and k explicitly. However, from Figure 9(b) where p = 3 and k = 1, it can be observed that the R1 population dwindles as they convert to either the P or M strategy. But by the tenth iteration, both M and P strategies appear to coexist (with an increase in agents with the P strategy in the later iterations), although complete convergence to a single strategy has not been achieved. By applying a higher cost of punishing agents (p = 6) compared to the cost of administering punishments (k = 1), the M population overtakes completely and quickly, as depicted in Figure 9(c).
@@
In the second experimental setup, the initial percentage of R1 agents increased from 25% to 37.5% (when compared to the first experimental setup), resulting in R1 agents being the largest group. Similar to the previous experiment, as portrayed in Figure 10(a), the R1 agent population dominates when the LLM is not provided with explicit punishment costs for both p and k. However, as shown in Figure 10(b), when punishment costs are set to p=3 and k=1, R1 agents change their behaviour, with all agents adopting the M strategy. However, when values for k and p are set to 6 and 1 respectively, unlike the first experiment, we observe that the R1 strategy is displaced by M and P strategies (Figure 10(c)). Since both M and P agents gain equal payoffs at each iteration, their strategy choices oscillate (i.e., without converging to a single strategy) due to agents’ random action selection decisions.
Both Figures 9 and 10 illustrate the effect of greater punishment costs (p) relative to the cost of punishing (k) on limiting the spread of the R1 strategy. Higher punishment costs incentivize agents to adopt cooperative strategies, leading to a complete shift away from other strategies, including Reluctant Cooperation, which aligns with the results of the B&R model. However, this outcome is influenced by the stochastic nature of the Fermi process (e.g., the oscillation observed in Figures 9(b) and 10(c)). Further systematic testing is required to statistically validate these findings and ensure a high-confidence assessment of strategy evolution.
@@ 6 Conclusions and Future Works
We have investigated whether the abstract mathematical evolution of cooperation studies conducted by B&R still holds in a more realistic simulation of a diner’s dilemma, where LLM agents make decisions and reason in natural language and adapt their strategies through the Fermi pairwise imitation mechanism. Our preliminary results indicate promising trends towards the evolution of cooperation given the explicit punishment values (i.e., the LLM is provided with explicit punishment costs for both p and k). However, though we observe the agents’ behaviours converge to the cooperative strategies (M & P) with punishment, it is subject to the random decision process implemented in the Fermi process. Moreover, longer iterations of the simulation will be necessary to investigate the results of Figures 9(b) and 10(c) where two cooperative strategies, M and P, are competing, with no agents left to defect. Therefore, additional systematic testing is necessary to confirm the results and validate the evolution of strategies with high confidence.
Additionally, throughout our experiments, we encountered several challenges, and as a part of solving those, we obtained insights that shaped our approach. Prompt engineering was one of the crucial steps, where overly complex and lengthy prompts led to inconsistent responses and hallucinated reasoning with LLMs, especially those with fewer parameters (70b in our case). Thus, we spent a considerable amount of time fine-tuning our prompts and testing them to obtain accurate results. Additionally, long-running simulations (around 15 minutes per iteration and 2.5 hours per simulation, 10 iterations in total) make large-scale experiments challenging, especially with the free-tier LLM request limits and the use of open-source LLMs. This underscores the need for further experiments with different LLM models, highlighting the areas for future improvements in scalability and robustness.
@@
Finally, in the future, we plan to systematically explore the long-term evolution of strategies over extended iterations, and different combinations of strategies in the population to solidify these preliminary findings and address the previously mentioned limitations.

### v3

@@ 5 Preliminary Results and Discussion
The dilemma decision accuracy reached 100%, meaning that the LLM correctly interpreted the natural language expression of the agent’s strategy to choose an action. Similarly, punishment accuracy achieved 100% across all strategies and conditions, indicating that agents correctly identified and punished defectors and non-punishers according to their strategies. Furthermore, this suggests that the influence of agent lifestyles was secondary. The LLM agents aligned their decisions with the behavioural strategy, showing no bias from additional factors such as lifestyles.
@@
The results of the first experimental setup (Fig. 9) show that without explicit punishment values, the Cooperative Punisher (P) agents have overtaken by the last iteration and the defecting agents (agents with R1) have survived until the ninth iteration, as shown in Figure 9(a). However, from Figure 9(b) where p = 3 and k = 1, it can be observed that the R1 population dwindles as they convert to either the P or M strategy, and by the tenth iteration, the Moralist strategy appears to be dominating, converging to a single strategy. By applying a higher cost of punishing agents (p = 6) compared to the cost of administering punishments (k = 1), the M population overtakes other strategies completely and quickly, as depicted in Figure 9(c). From all three above experiments, we can observe that with explicit punishment, the number of agents with the R1 strategy has diminished rapidly compared to the other variations.
@@
In the second experimental setup, the initial percentage of R1 agents increased from 25% to 37.5% (when compared to the first experimental setup), resulting in R1 agents being the largest group. As portrayed in Figure 10(a), now we can observe that the R1 agent population dominates when the LLM is not provided with explicit punishment costs for both p and k. However, as shown in Figure 10(b), when punishment costs are set to p=3 and k=1, R1 agents change their behaviour, with all agents adopting either M or P strategy, coexisting up until the fifth iteration and finally converging to the P strategy. Furthermore, when the value of p was increased to 6 with k=1, similar to the first experiment, we observe that the R1 strategy is displaced by M and P strategies (Figure 10(c)), and the population finally converges to Moralists.
In the first combination, where Moralists make up the majority of the population, LLM agents tend to shift toward cooperation primarily due to punishment and meta-enforcement strategies. Enforced punishments will reduce the utility of the non-cooperative agents, pushing R1 agents to adopt cooperative strategies. The emergence of the Moralist strategy ensures and sustains cooperation by punishing other agents who don’t punish defection, resulting in strategies that punish defection to remain in the population and continue penalizing defection. This trend is especially evident when many Reluctant-Cooperators are present, as seen in the 2nd combination experiments, where non-cooperation can be more profitable long term when the effect of punishment is not significant, echoing the findings by B&R [6]. It was only possible to allow Moralists to dominate the initially non-cooperative population, through increasing the cost of punishment (p). For this purpose, the agent behavioural strategy definitions used by the agents have to be clearly indicated to ensure the LLM agents choose the correct decisions and abide by their strategy behaviour.
Both Figures 9 and 10 illustrate the effect of greater punishment costs (p) relative to the cost of punishing (k) on limiting the spread of the R1 strategy. Higher punishment costs incentivize agents to adopt cooperative strategies, leading to a complete shift away from other strategies, including Reluctant Cooperation, which aligns with the results of the B&R model. However, this outcome is influenced by the stochastic nature of the Fermi process (e.g., the oscillation observed in Figures 10(b) and 10(c), iterations where there are only M and P strategies present in the population. In these iterations, both strategies gain equal payoffs at each iteration). Further systematic testing is required to statistically validate these findings and ensure a high-confidence assessment of strategy evolution.
@@ 6 Conclusions and Future Works
We have investigated whether the abstract mathematical evolution of cooperation studies conducted by B&R still holds in a more realistic simulation of a Diner’s Dilemma, where LLM agents make decisions and reason in natural language and adapt their strategies through the Fermi pairwise imitation mechanism. Our preliminary results indicate promising trends towards the evolution of cooperation given the explicit punishment values (i.e., the LLM is provided with explicit punishment costs for both p and k). Further, increased punishment costs caused populations with greater numbers of non-cooperative agents to converge to Moralist behaviours. However, though we observe the agents’ behaviours converge to the cooperative strategies (M & P) with punishment, it is subject to the random decision process implemented in the Fermi process. Moreover, longer iterations of the simulation, scaling up the number of agents, and additional systematic testing are necessary to confirm the results and validate the evolution of strategies with high confidence.
Additionally, throughout our experiments, we encountered several challenges, and as a part of solving those, we obtained insights that shaped our approach. Prompt engineering was one of the crucial steps, where overly complex and lengthy prompts led to inconsistent responses and hallucinated reasoning with LLMs, especially those with fewer parameters (70b in our case). Thus, we spent a considerable amount of time refining our prompts and testing them to obtain accurate results. Additionally, long-running simulations make large-scale experiments challenging, especially with the free-tier LLM request limits and the use of open-source LLMs. This underscores the need for further experiments with different LLM models, highlighting the areas for future improvements in scalability and robustness.
@@
Finally, in the future, we plan to systematically explore the long-term evolution of strategies over extended iterations and different combinations of strategies in the population to solidify these preliminary findings and address the previously mentioned limitations.